OpenStudy (anonymous):

I WILL GIVE YOU A MEDAL Use the formula for the sum of the first n terms of a geometric sequence to solve. Find the sum of the first 8 terms of the geometric sequence: -8, -16, -32, -64, -128, . . . . A. -2003 B. -2040 C. -2060 D. -2038

OpenStudy (anonymous):

is the answer B or C yes or no?

OpenStudy (amorfide):

use the formula (A(1-r^n))/1-r

OpenStudy (compassionate):

The n-th term of a geometric sequence with initial value 'a' and common ratio 'r' is given by: a (base 'n') = a r ^ (n-1) for e.g: Consider the following: 1, -3, 9, -27, 81, -243, ... The common ratio 'r' = (-3/1) = -3 OR = (81/-27) = -3

OpenStudy (compassionate):

The n-th term of a geometric sequence with initial value 'a' and common ratio 'r' is given by: a (base 'n') = a r ^ (n-1) for e.g: Consider the following: 1, -3, 9, -27, 81, -243, ... The common ratio 'r' = (-3/1) = -3 OR = (81/-27) = -3

OpenStudy (compassionate):

Plug in your terms...

OpenStudy (compassionate):

Are you aware how to use the formula?

OpenStudy (anonymous):

'a' TIMES (1-r^n)/(1-r) a = -8 r = 2 n = 8 what do i do from here

OpenStudy (anonymous):

-127/-1 ? or am i doing it wrong

OpenStudy (amorfide):

okay your formula is